# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v2

@@
The goal of influence maximization (IM) is to select a small set of seed nodes which maximizes the spread of influence on a network. In this work, we propose $\mathsf{BOPIM}$, a Bayesian Optimization (BO) algorithm for IM on temporal networks. The IM task is well-suited for a BO solution due to its expensive and complicated objective function. We propose a simple surrogate function to model the objective function and leverage Gaussian Process regression with shrinkage priors to fit the model. A major difficulty in combinatorial BO is constructing an appropriate covariance matrix. We overcome this challenge with a kernel function based on the amount of overlap in seed set neighbors, thus tailoring the solution to our IM application. This also allows us to use the Expected Improvement acquisition function to choose the next point to evaluate. In numerical experiments on real-world networks, we prove that $\mathsf{BOPIM}$ outperforms competing methods and yields comparable influence spreads to a gold-standard greedy algorithm while being as much as five times faster. We also demonstrate the proposed method’s ability to quantify uncertainty in optimal seed sets. To our knowledge, this is the first attempt to look at uncertainty in the seed sets for IM.
Keywords: Diffusion, Dynamic networks, Shrinkage priors, Uncertainty quantification
@@ 1 Introduction
Influence maximization (IM) is a canonical network science problem where the goal is to select a small set of seed nodes in order to maximize the influence spread on a graph (Kempe et al.,, 2003; Chen et al.,, 2009). Over the past two decades, IM has garnered significant research interest due to its relevance in online marketing campaigns (Domingos and Richardson,, 2001; Bharathi et al.,, 2007; Chen et al., 2010b, ), rumor spreading (Kempe et al.,, 2003; Murata and Koga,, 2018), public health campaigns (Wilder et al.,, 2017; Yadav et al.,, 2017; Wilder et al.,, 2018; Yadav et al.,, 2018), vaccine strategies (Holme,, 2004; Lee et al.,, 2012), and more. Traditionally, the problem assumes the graph to be static, i.e., nodes and edges are fixed. In many situations, however, this may be unrealistic. For example, consider a marketing campaign on X where a company promotes a new product. As users and followers change daily, an effective IM algorithm must account for this dynamism. Therefore, recent work has looked at the IM problem on temporal networks (Holme and Saramäki,, 2012) where edges change with time. Since this combinatorial optimization problem is NP-hard, heuristics must be employed. Some methods use a greedy algorithm and probabilistic approximation of the expected influence spread (Aggarwal et al.,, 2012; Osawa and Murata,, 2015; Erkol et al.,, 2020), while others eschew costly influence spread calculations by simply ranking nodes in order of importance (Michalski et al.,, 2014; Murata and Koga,, 2018; Michalski et al.,, 2020). We refer the interested reader to Yanchenko et al., 2024b for a review of temporal IM.
@@
In this work, we propose $\mathsf{BOPIM}$, Bayesian OPtimization for Influence Maximization, a novel BO algorithm for IM on temporal networks, which is one of the first solutions to the IM problem from a statistical angle. We construct an intuitive surrogate model for the influence spread function as well as a novel kernel function based on the similarity in neighbors of the seed set. We then leverage a popular global-local shrinkage prior and Expected Improvement acquisition function to select the next candidate point. In numerical experiments on real-world networks, $\mathsf{BOPIM}$ yields seed sets with influence spreads comparable to those of a greedy algorithm but at a fraction of the computational time. Additionally, our approach yields measures of uncertainty on the seed sets.
There have been many contributions to the combinatorial BO literature, including Baptista and Poloczek, (2018); Oh et al., (2019); Daulton et al., (2022); Papenmeier et al., (2023). These methods cannot be applied directly to our problem, however, due to the cardinality constraint on the number of node in the seed set. Our work primarily builds on Baptista and Poloczek, (2018); Garnett et al., (2010); Buathong et al., (2020). We leverage the ideas from Baptista and Poloczek, (2018) to construct the mean function for our Gaussian process, but this requires some modifications thanks to the cardinality constraint and input space size. Garnett et al., (2010); Buathong et al., (2020), on the other hand, address the challenges of constructing a covariance function for combinatorial optimization problems. Their approaches cannot be used directly either, however, because our input space is nodes in a graph, i.e., non-Euclidean. Though recently, there has been some work on BO for functions defined on graphs (Wan et al.,, 2024).
The layout of the paper is as follows. For the remainder of Section 1, we introduce notation and give a formal problem statement. Section 2 presents the main algorithm while Section 3 is devoted to numerical experiments. We share concluding thoughts and future directions in Section 4.
@@ 1.1 Notation and problem statement
Central to this problem is the influence diffusion mechanism. Some of the most popular choices are the Independent Cascade (IC) (Wang et al.,, 2012; Wen et al.,, 2017) and Linear Threshold (Chen et al., 2010a, ; Goyal et al.,, 2011) models. We adopt the Susceptible-Infected (SI) model from epidemiological studies (Osawa and Murata,, 2015; Murata and Koga,, 2018). In the SI model, a node is either in a susceptible (S) [^2] We use S to refer both to the susceptible state and seed set, but from context the usage should be clear. or infected (I) state. A node is infected if it was influenced at some point during the process, and susceptible otherwise. At time $t=1$, nodes in the seed set are set to state I, and all other nodes initialize to state S. At time $t$, a node in state I attempts to influence all of its neighbors. Specifically, if node $i$ is in state I, node $j$ is in state S, and there exists an edge between node $i$ and $j$ at time $t$, then node $j$ will be in state I at time $t+1$ with probability $\lambda$. The process concludes at time $T+1$ and the number of nodes in state I corresponds to the total influence spread. If we let $\sigma(S)$ be the expected number of influenced nodes at time $T+1$ for seed set $S$, the IM problem seeks the set of nodes of size $k$ that maximizes $\sigma(S)$ on $\mathcal{G}$, i.e.,
@@ 2 Bayesian Optimization
We now present the main contribution of this paper, the $\mathsf{BOPIM}$ algorithm, by addressing the three main steps of the BO schema: the objective function, statistical model and acquisition function.
@@ 2.1 Objective function
The first step in a BO algorithm is the initial evaluation of the objective function. Let $f:\mathcal{X}\to\mathbb{R}$ be some real-valued function defined on domain $\mathcal{X}$ that we wish to maximize. BO is particularly useful when $f(\cdot)$ is computationally expensive to evaluate, lacks special structure, e.g., concavity, and/or has unavailable first and second order derivatives. Then we evaluate $f(\bm{x}_{i})$ for $\bm{x}_{i}\in\mathcal{X}$ and $i=1,\dots,N_{0}$ for a user-defined $N_{0}$.
@@ 2.2 Statistical model
Next we need an adequate surrogate model for the objective function. The most popular approach uses Gaussian Process (GP) regression. First, let
$y_{i}=f(\bm{x}_{i})+\epsilon_{i},$
where $y_{i}$ is the observed influence spread, $f(\bm{x}_{i})$ is the objective function from Section 2.1 and $\epsilon_{i}\sim\mathsf{Normal}(0,\sigma^{2})$ is the noise term. Then our GP regression model is
$\bm{y}\sim\mathsf{Normal}(\mu({\bf X}),\sigma^{2}\gamma{\bf K}+\sigma^{2}{\bf I}_{N_{0}})$ (2)
where $\mu(\cdot)$ is the mean function and ${\bf K}$ is the covariance defined by a kernel function $\kappa(\cdot,\cdot)$ such that ${\bf K}_{ij}=\kappa(\bm{x}_{i},\bm{x}_{j})$. Additionally, $\bm{y}=(y_{1},\dots,y_{N_{0}})^{T}$, ${\bf X}$ is defined in the previous sub-section, and $\gamma>0$ is a kernel hyper-parameter. The two key components of the GP regression model are the mean function $\mu(\cdot)$ and kernel function $\kappa(\cdot,\cdot)$, so we consider each of these in turn.
@@ 2.2.1 Mean function
First, we construct a mean function, $\mu(\cdot)$ parameterized by $\bm{\beta}$, to serve as a surrogate for $f(\cdot)$. Motivated by Baptista and Poloczek, (2018), we assume that the surrogate model is linear in its parameters, i.e.,
$\mu(\bm{x}_{i})=\sum_{j=1}^{n}x_{ij}\beta_{j}=\bm{x}_{i}\bm{\beta}$ (3)
$\beta_{j}|\lambda_{j}^{2},\tau^{2},\sigma^{2}$ $\sim\mathsf{Normal}(0,\lambda_{j}^{2}\tau^{2}\sigma^{2})$
$\lambda_{j}^{2},\tau^{2}$ $\sim\mathsf{C}^{+}(0,1)$ (4)
where $j=1,\dots,n$ and $\mathsf{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Thus, $\tau^{2}$ represents the global shrinkage and $\lambda_{j}^{2}$ controls the local shrinkage of each coefficient. For the Gibbs sampler, we leverage an auxiliary variable formulation (Makalic and Schmidt,, 2015). Please see the Supplemental Materials for complete MCMC details. We also considered the R2D2 prior (Zhang et al.,, 2022; Yanchenko et al., 2024a, ) and Dirichlet-Laplace (Bhattacharya et al.,, 2015), but the results were similar, so we focus on the HS prior.
2.2.2 Coefficient interpretation
@@ 2.2.3 Kernel function
To complete our GP regression model, we must also specify the covariance, ${\bf K}$, defined by the kernel function, $\kappa(\cdot,\cdot)$. If $\bm{x}_{i},\bm{x}_{j}\in\mathbb{R}^{p}$, then it is straightforward to compute their variance using, e.g., a Gaussian or Matern kernel. In our setting, however, $\bm{x}_{i},\bm{x}_{j}\in\{0,1\}^{n}$ and correspond to nodes in a graph, so it is not immediately obvious how to choose a kernel.
@@
With kernel function in hand, we can easily construct our covariance matrix. Indeed,
${\bf K}_{ij}=\kappa(\bm{x}_{i},\bm{x}_{j})=\begin{cases}JC(\mathcal{S}_{i},\mathcal{S}_{j})&i\neq j\\ 1+\delta&i=j\end{cases}$ (5)
where $\delta>0$ is some small constant to ensure positive-definiteness of our covariance matrix. This definition satisfies the desideratum that points close in the input space are more strongly correlated where we have defined two seed sets to be “close” if they have many neighbors in common at time $t=1$.

### v4

@@
The goal of influence maximization (IM) is to select a small set of seed nodes which maximizes the spread of influence on a network. In this work, we propose $\mathsf{BOPIM}$, a Bayesian Optimization (BO) algorithm for IM on temporal networks. The IM task is well-suited for a BO solution due to its expensive and complicated objective function. There are at least two key challenges, however, that must be overcome, primarily due to the inputs coming from a cardinality-constrained, non-Euclidean, combinatorial space. The first is constructing the kernel function for the Gaussian Process regression. We propose two kernels, one based on the Hamming distance between seed sets and the other leveraging the Jaccard coefficient between node’s neighbors. The second challenge is the acquisition function. For this, we use the Expected Improvement function, suitably adjusting for noise in the observations, and optimize it using a greedy algorithm to account for the cardinality constraint. In numerical experiments on real-world networks, we prove that $\mathsf{BOPIM}$ outperforms competing methods and yields comparable influence spreads to a gold-standard greedy algorithm while being as much as ten times faster. In addition, we find that the Hamming kernel performs favorably compared to the Jaccard kernel in nearly all settings, a somewhat surprising result as the former does not explicitly account for the graph structure. Finally, we demonstrate two ways that the proposed method can quantify uncertainty in optimal seed sets. To our knowledge, this is the first attempt to look at uncertainty in the seed sets for IM.
Keywords: Diffusion, Dynamic networks, Hamming distance, Jaccard coefficient, Shrinkage priors
@@ 1 Introduction
Influence maximization (IM) is a canonical network science problem where the goal is to select a small set of seed nodes in order to maximize the influence spread on a graph (Kempe et al.,, 2003; Chen et al.,, 2009). Over the past two decades, IM has garnered significant research interest due to its relevance in online marketing campaigns (Domingos and Richardson,, 2001; Bharathi et al.,, 2007; Chen et al., 2010b, ), rumor spreading (Kempe et al.,, 2003; Murata and Koga,, 2018), public health campaigns (Wilder et al.,, 2017; Yadav et al.,, 2017; Wilder et al.,, 2018; Yadav et al.,, 2018), vaccine strategies (Holme,, 2004; Lee et al.,, 2012), and more. Traditionally, the problem assumes the graph to be static, i.e., nodes and edges are fixed. In many situations, however, this may be unrealistic. For example, consider a marketing campaign on X where a company promotes a new product. As users and followers change daily, an effective IM algorithm must account for this dynamism. Therefore, recent work has looked at the IM problem on temporal networks (Holme and Saramäki,, 2012) where edges change with time. Since this combinatorial optimization problem is NP-hard, heuristics must be employed. Some methods use a greedy algorithm and probabilistic approximation of the expected influence spread (Aggarwal et al.,, 2012; Osawa and Murata,, 2015; Erkol et al.,, 2020), while others eschew costly influence spread calculations by simply ranking nodes in order of importance (Michalski et al.,, 2014; Murata and Koga,, 2018; Michalski et al.,, 2020). We refer the interested reader to Yanchenko et al., (2024) for a review of temporal IM.
@@
In this work, we propose $\mathsf{BOPIM}$, Bayesian OPtimization for Influence Maximization, a BO algorithm for IM on temporal networks. The key challenges in applying a BO framework to our problem are in constructing the kernel function and optimizing the acquisition function. To address these, we propose two kernels, one based on the Hamming distance between seed sets, and the other on the similarity in node’s neighbors. We then leverage the Expected Improvement acquisition function to select the next candidate point, appropriately accounting for the noise in our observations. We optimize this function using a greedy algorithm which shares many similarities to the Trust Region framework (e.g., Eriksson et al.,, 2019; Wan et al.,, 2021; Papenmeier et al.,, 2023), but also accounts for the cardinality constraint. In numerical experiments on real-world networks, $\mathsf{BOPIM}$ yields seed sets with influence spreads comparable to those of a “gold-standard” greedy algorithm but at a fraction of the computational time.
There are two primary contributions of this work. First, to the knowledge of the author, this is that first time that a BO framework has been applied to the IM problem. The key challenges stem from the inputs having a cardinality constraint, while also corresponding to non-Euclidean space, i.e., nodes on a graph. There have been several contributions to the combinatorial BO literature, including Garnett et al., (2010); Baptista and Poloczek, (2018); Oh et al., (2019); Ru et al., (2020); Buathong et al., (2020); Daulton et al., (2022); Papenmeier et al., (2023). The proposed method is inspired by these papers, most notably in the construction of the kernel, and optimization of the acquisition function. The cardinality constraint, however, means that these methods cannot easily be applied directly. For example, we show in detail why $\mathsf{COMBO}$ (Oh et al.,, 2019) cannot be used with such a constraint. Thus, this work addresses the various challenges of adapting the BO framework to the IM problem.
Second, the BO algorithm naturally provides measures of uncertainty on the optimal seed sets. While there are many methods developed for IM, there has been little emphasis on the uncertainty in the results. The proposed method not only provides uncertainty quantification of the total influence spread on the network, but also a richer understanding of a node’s importance in the optimal seed set. Indeed, $\mathsf{BOPIM}$ can inform us if there are many seed sets which yield similar influence spreads, and/or how confident we are that a node should be in the optimal seed set, as opposed to a simple binary result on its inclusion/exclusion. We discuss two different ways that $\mathsf{BOPIM}$ can help answer these questions.
The layout of the paper is as follows. For the remainder of Section 1, we introduce notation and give a formal problem statement. Section 2 presents the main algorithm while Section 3 is devoted to numerical experiments. In Section 4 we discuss uncertainty quantification, and share concluding thoughts and future directions in Section 5.
@@ 1.1 Notation and problem statement
We first must choose an influence diffusion mechanism, though the proposed method can easily be adapted to any diffusion mechanism. Some of the most popular choices are the Independent Cascade (IC) (Wang et al.,, 2012; Wen et al.,, 2017) and Linear Threshold (Chen et al., 2010a, ; Goyal et al.,, 2011) models. We adopt the Susceptible-Infected (SI) model from epidemiological studies (Osawa and Murata,, 2015; Murata and Koga,, 2018). In the SI model, a node is either in a susceptible ($\mathscr{S}$) or infected ($\mathscr{I}$) state. A node is infected if it was influenced at some point during the process, and susceptible otherwise. At time $t=1$, nodes in the seed set are set to state $\mathscr{I}$, and all other nodes initialize to state $\mathscr{S}$. At time $t$, a node in state $\mathscr{I}$ attempts to influence all of its neighbors. Specifically, if node $i$ is in state $\mathscr{I}$, node $j$ is in state $\mathscr{S}$, and there exists an edge between node $i$ and $j$ at time $t$, then node $j$ will be in state $\mathscr{I}$ at time $t+1$ with probability $\lambda$. The process concludes at time $T+1$ and the number of nodes in state $\mathscr{I}$ corresponds to the total influence spread. If we let $\sigma(S)$ be the expected number of influenced nodes at time $T+1$ for seed set $S$, the IM problem seeks the set of nodes of size $k$ that maximizes $\sigma(S)$ on $\mathcal{G}$, i.e.,
@@ 2 Bayesian Optimization
In this section, we present the main contribution of this paper, the $\mathsf{BOPIM}$ algorithm, particularly emphasizing the unique challenges posed by the temporal IM problem.
@@ 2.1 Objective function
$y=f(\bm{x})+\epsilon,$
where $y$ is the observed influence spread, $f(\bm{x})$ is the objective function and $\epsilon\sim\mathsf{Normal}(0,\sigma^{2})$ is the noise term.
2.2 Initial evaluations
We initially obtain realizations of the spreading process for $N_{0}$ seed sets. We remark that nodes with high degrees on the temporally aggregated network, $\tilde{G}=\cup_{t}G_{t}$, are good candidates for the optimal seed set. Therefore, instead of randomly sampling $\bm{x}_{i}$ from $\mathcal{X}$, we sample nodes proportionally to their degree. Mathematically, this means sampling $k$ times without replacement from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k}d_{k}$ where $d_{j}$ is the degree of node $j$ on $\tilde{G}$ for $Z\in\{1,\dots,n\}$. This sampling scheme ensures that the model better fits the objective function near the expected optimal seed set (see Supplemental Materials for empirical evidence). With $\bm{x}_{1},\dots,\bm{x}_{N_{0}}$ sampled, we define ${\bf X}=(\bm{x}_{1}^{T},\dots,\bm{x}_{N_{0}}^{T})^{T}\in\{0,1\}^{N_{0}\times n}$ and obtain $\bm{y}=(y_{1},\dots,y_{N_{0}})^{T}$ via MC simulations. Note that if identical seed sets are sampled, we discard one and re-sample to ensure $N_{0}$ unique initial seed sets. In the Supplemental Materials, we also discuss an alternative initial sampling scheme.
@@ 2.3 Statistical model
Next, we need an adequate surrogate model for the objective function. The most popular approach uses Gaussian Process (GP) regression. As the influence spreads are observed with noise, we assume the following GP regression model:
$\bm{y}\sim\mathsf{Normal}(\mu({\bf X}),\sigma^{2}\gamma{\bf K}+\sigma^{2}{\bf I}_{N_{0}}),$ (2)
where $\mu(\cdot)$ is the mean function and ${\bf K}$ is the correlation matrix defined by a kernel function $\kappa(\cdot,\cdot)$ such that $K_{ij}=\kappa(\bm{x}_{i},\bm{x}_{j})$. Moreover, $\gamma>0$ is a kernel hyper-parameter, and $\bm{y}$ and ${\bf X}$ are defined in the previous sub-section. The two key components of the GP regression model are the mean function $\mu(\cdot)$ and kernel function $\kappa(\cdot,\cdot)$, so we consider each in turn.
@@ 2.3.1 Mean function
First, we must specify the mean function $\mu(\cdot)$ of our GP regression model. As is typical, we consider an intercept-only model, i.e.,
$\mu(\bm{x})=\beta_{0}{\bf 1}_{N_{0}}$
for all $\bm{x}$, where $\beta_{0}$ is the intercept and ${\bf 1}_{n}$ is a vector of ones of length $n$. In Section 4, we introduce another mean function to allow for better uncertainty quantification of the optimal seed set.
@@ 2.3.2 Kernel function
To complete our GP regression model, we must also specify the correlation matrix, ${\bf K}$, defined by the kernel function, $\kappa(\cdot,\cdot)$. If $\bm{x}_{i},\bm{x}_{j}\in\mathbb{R}^{p}$, then it is straightforward to compute their covariance using, e.g., a Gaussian or Matern kernel. In our setting, however, $\bm{x}_{i},\bm{x}_{j}\in\{0,1\}^{n}$ and correspond to nodes in a graph, so it is not immediately obvious how to choose a kernel. We discuss several possible choices.
COMBO
Recently, $\mathsf{COMBO}$ was proposed as an efficient and scalable method for combinatorial Bayesian optimization (Oh et al.,, 2019). This paper proposes a novel kernel function for combinatorial search spaces based on the idea of a combinatorial graph, where each vertex [^3] To attempt to avoid confusion between the combinatorial graph and the original graph where the information diffusion occurs, we will refer to the former as being made up of vertices, and the latter as possessing nodes. corresponds to a possible input to the objective function, and there is an edge between vertices if the inputs differ by only a single variable. The authors construct such a graph using the graph Cartesian product and derive a kernel based on the Graph Fourier Transform.
In the context of IM, each vertex in the combinatorial graph would correspond to a possible seed set and there would be an edge between vertices if the corresponding seed sets only differed by one node. While we can construct such a graph, the cardinality constraint imposed on the seed sets means that we cannot construct the graph using the graph Cartesian product. Therefore, it is not straightforward to apply the ideas of $\mathsf{COMBO}$ to the IM problem. Please see the Supplemental Materials for further discussion of why the method is not applicable. We consider extending $\mathsf{COMBO}$ to constrained input spaces an interesting avenue of future work.
Hamming Distance
In Oh et al., (2019), the authors show that the shortest distance between nodes on the combinatorial graph is equivalent to the Hamming distance. While we showed that $\mathsf{COMBO}$ cannot be used in our problem, we can borrow these ideas to construct a kernel function based on the Hamming distance. Specifically, let $\bm{x}_{i}$ and $\bm{x}_{j}$ be two seed sets. Then the covariance matrix is defined by
$K_{ij}=\kappa(\bm{x}_{i},\bm{x}_{j})=\begin{cases}1-\frac{1}{2k}d_{H}(\bm{x}_{i},\bm{x}_{j})&i\neq j\\ 1+\delta&i=j\end{cases}$ (3)
where $\delta>0$ is some small constant to ensure positive-definiteness of our covariance matrix. In all experiments, we set $\delta=0.01$. Additionally, $d_{H}(\bm{u},\bm{v})$ is the Hamming distance, which counts the number of entries of between vectors $\bm{u}$ and $\bm{v}$ which are not equal, i.e.,
$d_{H}(\bm{u},\bm{v})=\sum_{i=1}^{n}\mathbb{I}(u_{i}\neq v_{i})$
where $\mathbb{I}(\cdot)$ is the indicator function. Because we are restricted to $k$ nodes in the seed set, if $k\leq n/2$, then $0\leq d_{H}\bm{x}_{i},\bm{x}_{j})\leq 2k$, which is why we divide by $2k$ instead of $n$. This kernel is in a similar spirit to those of $\mathsf{CASMOPOLITAN}$ (Wan et al.,, 2021), $\mathsf{Bounce}$ (Papenmeier et al.,, 2023) and $\mathsf{CoCaBO}$ (Ru et al.,, 2020). Please see the Supplemental Materials for a proof that this kernel is positive semi-definite.
Jaccard Coefficient
While the Hamming distance is a sensible choice for a kernel metric, a potential drawback is that it does not account for the structure of the graph. Instead, it simply enumerates the number of nodes shared between seed sets. But since the graph structure, node neighbors, etc. affect the influence spread, our kernel function should seemingly also account for these properties.
@@
With kernel function in hand, we can easily construct our covariance matrix:
$K_{ij}=\kappa(\bm{x}_{i},\bm{x}_{j})=\begin{cases}JC(\mathcal{S}_{i},\mathcal{S}_{j})&i\neq j\\ 1+\delta&i=j\end{cases}$ (4)
where again, $\delta>0$ is some small constant ($\delta=0.01$). This definition satisfies the desideratum that points close in the input space are more strongly correlated where we have defined two seed sets to be “close” if they have many neighbors in common at time $t=1$.
The use of the JC in such context enjoys some precedence. First, the JC has been used for link prediction in graphs, most notably in (Liben-Nowell and Kleinberg,, 2003). Additionally, Gosnell and Evangelou, (2024) use the Tanimoto index, another name for the JC, to construct a GP kernel with drug discovery applications. More generally, the JC has received significant attention in chemoinformatics (Rogers and Tanimoto,, 1960; Gower,, 1971; Todeschini et al.,, 2012; Moss and Griffiths,, 2020; Gosnell and Evangelou,, 2024). Finally, the JC (Tanimoto index) has been shown to yield a positive semi-definite correlation matrix, making it a valid choice for our kernel function (Gower,, 1971; Bouchard et al.,, 2013).